# MASt3R-Fusion: Integrating Feed-Forward Visual Model with IMU, GNSS for High-Functionality SLAM

### v2

@@ Abstract
I Introduction
Visual SLAM is a widely utilized technique in applications such as robotics, autonomous driving, and extended reality (XR)[1]. SLAM systems generally require comprehensive functionality to support diverse spatial tasks. Compared with 3D reconstruction, visual SLAM places a stronger emphasis on real-time pose estimation and scene structure perception. In contrast to visual odometry, SLAM systems aim to provide a more complete understanding of the environment while simultaneously maintaining long-term global consistency.
Although classical visual SLAM frameworks have achieved notable successes, their practical usability remains limited due to several critical factors. First, traditional visual SLAM methods based on sparse features or photometric consistency often fail to achieve complete spatial perception, particularly in low-texture environments. Second, the system performance tends to degrade significantly under challenging visual conditions. It also remains an open problem how to achieve life-long consistency across time periods and under varying illumination conditions.
To overcome these limitations, the integration of deep learning/foundation models has emerged as an effective solution[2]. In particular, recent feed-forward neural network-based approaches for visual geometry estimation (e.g., DUSt3R[3], MASt3R[4], VGGT[5]), have attracted growing attention. These methods encode the images and decode the latents to directly recover 2D-to-3D pointmaps and other dense information. By leveraging large-scale data to learn spatial priors, they alleviate many of the degeneracies encountered in traditional multi-view stereo (MVS) and SLAM systems, which greatly inspire recent SLAM implementations[6, 7].
@@ I Introduction
Despite significant progress, vision-based methods still suffer from inherent limitations, especially regarding scale consistency and performance degradation in visually deprived environments. To address this, a promising direction is to integrate these novel visual paradigms with complementary sensors, such as inertial measurement units (IMUs) and global navigation satellite system (GNSS), thus to exploit the respective strengths of traditional multi-sensor fusion and learning-based spatial priors.
@@
    We propose a framework to fuse feed-forward pointmap regression with multi-sensor information in a tight way, bridging Sim(3)-based visual alignment constraints and metric-scale SE(3) states and factors.
@@
    Based on the feed-forward model, we develop a real-time visual-inertial SLAM system that supports metric-scale pose estimation and dense perception.
@@
II Related Work
II-A Feed-Forward Pointmap Regression
@@
Building upon DUST3R, MASt3R[4] introduces several key innovations that address critical bottlenecks, including a descriptor head for refined matching and metric scale inference. Recent improvements to the feed-forward model paradigm include multi-view extensions[8, 5], dynamic scenes[9], permutation equivariance[10], prior information assistance[11], and so on.
II-B Deep Visual SLAM
For visual SLAM systems, deep learning has introduced transformative changes beyond the classical implementations[12]. These advancements span both modular improvements, such as enhanced feature matching[13, 14], depth estimation[15, 16], and spatial representations including object-level SLAM[17, 18], renderable neural fields[19], and 3D Gaussian Splatting (3DGS)[20]. In addition, end-to-end approaches have also been explored, where images are directly used as input to estimate outputs such as camera poses and depth maps. Typically, DROID-SLAM[21] achieves an end-to-end SLAM framework by combining dense optical flow with a differentiable dense bundle adjustment (DBA) module. Notably, it retains probabilistic geometric estimation, resulting in high accuracy and robustness.
@@ II-B Deep Visual SLAM
II-C Multi-Sensor-Aided Visual SLAM
Incorporating multi-sensor information into visual SLAM systems is a common approach to enhance system functionality. A representative example is the integration of visual and inertial data, which enables high-precision, metric-scale pose estimation through filter- or optimization-based information fusion[23, 24]. More advanced methods include photometric-based measurement[25], delayed marginalization[26], and pose-only modeling[27].
@@ II-C Multi-Sensor-Aided Visual SLAM
Apart from IMU, some methods integrate GNSS information for global functionality[33, 34, 35]. However, it remains a challenge how to effectively integrate visual(/visual-inertial) inforamtion with GNSS for global optimal estimation, as existing methods either focus solely on real-time navigation or rely on pose graph optimization that loses a large amount of visual/inertial information.
@@ III System Overview
In the real-time SLAM stage, input images are processed through a feed-forward neural network to recover the pointmaps and achieve dense matching. Pose tracking is performed by aligning pointmaps between consecutive frames. A sliding-window factor graph is maintained over a fixed number of keyframes. Sim(3)-based visual constraints among covisible frames are established, while IMU data are tightly incorporated. Probabilistic marginalization is applied to maximize information utilization, which enables high-accuracy pose estimation and metric-scale structure recovery with low-drift performance.
@@
IV Visual Measurement based on Feed-Forward Model
@@
IV-A Two-View Feed-Forward Model
Our work uses MASt3R, a typical two-view feed-forward model to process the images, which facilitates pointmap regression and dense data association. The pointmap regression pipeline is depicted in Fig. 3. Firstly, images are encoded into feature tokens following
@@ IV-A Two-View Feed-Forward Model
For an image pair (denoted as $\mathbf{I}_{i}$ and $\mathbf{I}_{j}$), their tokens are decoded to produce the pointmaps and descriptors
@@
Thanks to large-scale training and the learning capability of transformers, the pointmap regression paradigm has been demonstrated to be highly effective. Essentially, the model integrates 3D spatial geometric priors with two-view stereo to produce reasonable structure information in various conditions.
As the regressed pointmaps are described in the unified reference frame $i$, dense matching based on ray proximity can be achieved through the following optimization
@@
where $\mathbf{u}^{i}_{j}$ with $(H\times W,2)$ shape denotes the 2D coordinates on image $i$ associated with the grid points on image $j$, $\left[\cdot\right]$” denotes retrieving the value at the corresponding 2D index on the pointmap. Here, bilinear interpolation is utilized for the retrieval, and the gradient map of $\mathbf{X}^{ij}_{i}$ is used for efficient optimization. After the optimization based on ray proximity, awareness of the 3D structure can help eliminate dynamic points (with large depth residuals).
As pointed by[4], geometry-based matching can be further improved by taking into account the feature-based refinement, which utilizes the descriptor maps following
@@
where $d(\cdot,\cdot)$ denotes the dot-product. As the the channel number of $\mathbf{D}$ is large and the descriptor lacks smoothness, a neighborhood searching (taking $\hat{\mathbf{u}}^{i}_{j}$ as initials) rather than optimization is utilized to obtain $(\hat{\mathbf{u}}^{i}_{j})^{\prime}$.
To achieve sub-pixel data association, we asymmetrically upsample $\mathbf{D}^{ij}_{j}$ through bilinear interpolation
@@
then perform the refinement around $(\hat{\mathbf{u}}^{i}_{j})^{\prime}$, following
@@
IV-B Visual Constraint based on Pointmap Alignment
Considering multiple frames of images, we maintain both the pointmaps and their associated poses, represented as
@@ IV-B Visual Constraint based on Pointmap Alignment
is the camera-to-world Sim(3) transformation. Here, $\mathbf{S}_{i}$ first scales the points to a unified scale and then transforms them into the world coordinate frame.
The self-referenced pointmap $\mathbf{X}_{i}$ is initialized using the two-view feed-forward results from Eq. (2), following
@@
Specifically, assuming known camera intrinsics, the 3D points can be constrained to lie along their corresponding rays
$\mathbf{X}_{i}\leftarrow\pi\left(\mathbf{u}_{i},\left(\mathbf{X}_{i}\right)_{z}\right)$ (10)
where $\mathbf{u}_{i}$ are the grid-like pixel coordinates, $\left(\mathbf{X}_{i}\right)_{z}$ is the depth vector, $\pi(\cdot)$ is the camera projection model, whose inverse operation backprojects 2D pixels to 3D points.
Although feed-forward reconstruction is considered advantageous in that it obviates the requirement for camera parameters, in practice we found that camera distortions can introduce considerable instability. Therefore, we consider the case where the camera parameters are known.
Figure 5: Illustration different forms of visual constraints. Left: bundle adjustment, in which landmarks and camera poses are jointly optimized with little prior knowledge. Right: Pointmap alignment, in which camera poses and pointmap scales are optimized, built upon the knowledge of the scaleless 3D structure.
Based on the dense association $\mathbf{u}^{i}_{j}$ produced by two-view matching, we can align two pointmaps and construct the following residuals
@@
where $\circ$” denotes the transforming operation. The residuals in the upper part resemble the reprojection error with known depth and play a dominant role, whereas the residuals in the lower part are primarily used to handle pure rotation cases.
The inter-frame transformation can be directly estimated based on the above pointmap alignment residuals
@@
where bi-directional projections are considered, $\mathbf{Q}_{ij}$, $\mathbf{Q}_{ji}$ are the weighting factors which consider network-outputted confidence and Huber robust kernel function. This is used for pose tracking of a new frame.
It is noted that, the residual computation and the optimization process are dense due to the large number of pixels. Therefore, both the residuals and the Hessian information for each projection are computed on GPU. To be specific, the Hessian inforamtion is computed following
@@
where $\bm{\eta}^{i}_{j}\in\mathbb{R}^{7}$ is the Lie algebra, $\mathbf{J}^{r}_{ij}$ is the $\mathbf{r}_{ij}$-to-$\mathbf{S}^{i}_{j}$ Jacobian. After GPU processing, the Hessian-form inforamtion $(\mathbf{H}_{ij},\mathbf{v}_{ij})$ are in compact $(7,7)$ and $(7,1)$ shapes and can be processed efficiently on CPU.
When multiple frames within a neighborhood are considered, residuals can be constructed for multiple image pairs, thereby enabling joint estimation of multi-frame poses and scales
@@
where $\mathcal{E}$ is the set of valid two-view projections. During the optimization, local-to-global Jacobians $\mathbf{J}^{ij}_{(i,j)}\in\mathbb{R}^{7\times 14}$ are utilized.
In contrast to BA commonly used in vision-based methods that jointly estimate point depths and camera poses, the above-described visual measurement model does not include point depths, and instead builds relatively independent constraints between pairwise images. This greatly simplifies the formulation of the problem, which is made possible by the strong capability of the visual foundation model to provide accurate estimates of 3D scene structure (up to a certain scale).
For comparison, we revisit the widely used BA technique and the re-prejection error formula, as shown below
$\mathbf{r}_{ij}\left(\mathbf{T}^{i}_{j},\mathbf{d}_{j}\right)=\left[\begin{matrix}\mathbf{u}^{i}_{j}-\pi\left(\mathbf{T}^{i}_{j}\circ\pi^{-1}\left(\mathbf{u}_{j},\mathbf{d}_{j}\right)\right)\end{matrix}\right]$ (16)
$\left\{\hat{\mathbf{T}},\hat{\mathbf{d}}\right\}=\arg\min_{\left\{\mathbf{T},\mathbf{d}\right\}}\sum_{i,j\in\mathcal{K}}\left\|\mathbf{r}_{ij}\left(\mathbf{T}_{i},\mathbf{T}_{j},\mathbf{d}_{j}\right)\right\|^{2}_{\mathbf{Q}_{ij}}$ (17)
where $\mathbf{d}_{j}$ are the depths of landmarks in the source frame $j$, $\mathbf{u}_{j}$ are the pixel coordinates.
It can be seen that when BA constructs visual constraints, it relies on the parameterization of a large number of point depths $\mathbf{d}_{j}$; this greatly increases the scale of the problem and needs careful initialization for good convergence. For better efficiency, a two-step optimization can be adopted by eliminating landmarks first through Schur complement and then constructing pose constraints[36]. However, this process involves all frames related to $\mathbf{d}_{j}$, making the construction of visual constraints less flexible. These issues no longer exist in pointmap-alignment-based visual constraints, which can be easily modeled as pairwise residuals.
It is noted that, this advantage relies on the assumption that the pointmap structure is accurate and scale is the only uncertain factor. Thanks to the spatial awareness of the feed-forward model, this assumption generally holds. Yet, under specific conditions, significant errors may be introduced—particularly when the points appear very far in image $j$ but very close in image $i$. In such cases, the depth uncertainty of the points in image $j$ will have a pronounced impact (large disparities) in image $i$. This issue is not obvious in small-scale scenes, but becomes significant in large-scale outdoor environments, typically for the long-time forward-moving scenario.
To mitigate the influence of this uncertainty, we simply apply a mask to downweight the residuals in the above mentioned projection process, as shown below
@@
V Real-Time SLAM
Based on the above paradigm, visual alignment constraints operate independently on a frame-to-frame basis, involving only a very compact set of associated parameters. Beyond the shared common parameters, there exist no functional dependencies among distinct visual constraints. This characteristic enables highly flexible and modular construction of factor graphs. Nevertheless, several issues need to be addressed:

### v3

@@ Abstract
To overcome these limitations, the integration of deep learning/foundation models has emerged as an effective solution[1]. In particular, recent feed-forward neural network-based approaches for visual geometry estimation (e.g., DUSt3R[2], MASt3R[3], VGGT[4]), have attracted growing attention. These methods encode the images and decode the latents to directly recover 2D-to-3D pointmaps and other geometric information. By leveraging large-scale data to learn the spatial priors, they alleviate many of the degeneracies encountered in traditional multi-view stereo (MVS) and SLAM systems, which greatly inspire recent SLAM implementations[5, 6].
Despite significant progress, vision-based methods still suffer from inherent limitations, especially regarding scale consistency and performance degradation in visually deprived environments. To address this, a promising direction is to integrate the novel visual paradigms with complementary sensors, such as inertial measurement units (IMUs) and global navigation satellite system (GNSS), thus to exploit the respective strengths of traditional multi-sensor fusion and learning-based spatial priors. The integration of multi-sensor information in visual SLAM systems has demonstrated remarkable effectiveness[7, 8, 9]; However, how to flexibly apply this approach to emerging vision-based methods remains to be explored.
@@
    We propose a framework to fuse feed-forward point-map regression paradigm with multi-sensor information in a tight way, bridging Sim(3)-based visual alignment constraints with metric-scale SE(3) states and factors.
@@
    Based on the feed-forward model, we develop a real-time visual-inertial SLAM system that supports metric-scale pose estimation and dense scene structure perception.
@@
I Related Work
I-A Feed-Forward Pointmap Regression
@@
Building upon DUST3R, MASt3R[3] introduces several key innovations that address critical bottlenecks, including a descriptor head for refined matching and the ability of metric scale inference. Recent improvements to the feed-forward model paradigm include multi-view extensions[10, 4], dynamic scenes[11], permutation equivariance[12], prior information assistance[13], and so on.
I-B Deep Visual SLAM
For visual SLAM systems, deep learning has introduced transformative changes beyond the classical implementations[14]. These advancements span both modular improvements, such as enhanced feature matching[15, 16], image retrieval[17], depth estimation[18, 19], and spatial representations including object-level SLAM[20, 21], renderable neural fields[22], and 3D Gaussian Splatting (3DGS)[23]. In addition, end-to-end approaches have also been explored, where images are directly used as input to estimate outputs such as camera poses and depth maps. Typically, DROID-SLAM[24] achieves an end-to-end SLAM framework by combining dense optical flow with a differentiable dense bundle adjustment (DBA) module. Notably, it retains probabilistic geometric estimation, resulting in high accuracy and robustness.
@@ I-B Deep Visual SLAM
I-C Multi-Sensor-Aided Visual SLAM
Incorporating multi-sensor information into visual SLAM systems is a common approach to enhance system functionality. A representative example is the integration of visual and inertial data, which enables high-precision, metric-scale pose estimation through filter- or optimization-based information fusion[7, 26]. More advanced methods include photometric-based measurement[27], specialized camera model[28, 29], delayed marginalization[30], and pose-only modeling[31].
@@ I-C Multi-Sensor-Aided Visual SLAM
Apart from IMU, some methods integrate GNSS information for geo-referenced functionality[9, 8, 37]. However, it remains a challenge how to effectively integrate visual(/-inertial) inforamtion with GNSS for global optimal estimation, as existing methods either focus solely on real-time navigation or rely on pose graph optimization that loses a large amount of visual/inertial information.
@@ II System Overview
In the real-time SLAM stage, input images are processed through a feed-forward neural network to recover the pointmaps and achieve further dense matching. Pose tracking is performed by aligning pointmaps between consecutive frames. A sliding-window factor graph is maintained over a fixed number of keyframes. Sim(3)-based visual constraints among co-visible frames are established, while IMU data are tightly incorporated. Probabilistic marginalization is applied to maximize information utilization, which enables high-accuracy pose estimation and metric-scale structure recovery with low-drift performance.
@@
III Visual Measurement based on Feed-Forward Model
@@
III-A Two-View Feed-Forward Model
Our work uses MASt3R, a typical two-view feed-forward model to process the images, which facilitates pointmap regression and dense data association. The pipeline of pointmap regression is depicted in Fig. 3. Firstly, images are encoded into feature tokens following:
@@ III-A Two-View Feed-Forward Model
For an image pair (denoted as $\mathbf{I}_{i}$ and $\mathbf{I}_{j}$), their tokens are decoded to produce the pointmaps and descriptors:
@@
Benefiting from large-scale training and the learning capability of transformers, the pointmap regression paradigm has been demonstrated to be highly effective. Essentially, the model integrates 3D spatial geometric priors with two-view stereo to produce reliable structure information under diverse conditions.
As the regressed pointmaps are described in the unified reference frame $i$, dense matching based on ray proximity can be achieved through the following optimization:
@@
where $\mathbf{u}^{i}_{j}$ with $(H\times W,2)$ shape denotes the 2D coordinates on image $i$ associated with the pixel grids on image $j$, $\left[\cdot\right]$” denotes retrieving the value at the corresponding 2D index of the pointmap. Here, bilinear interpolation is employed for the retrieval, and the gradient map of $\mathbf{X}^{ij}_{i}$ is used for efficient optimization. After the optimization based on ray proximity, point correspondences with large depth residuals are masked as invalid, which helps eliminate dynamic objects with the awareness of the 3D structure provided by the feed-forward model.
As pointed by MASt3R[3], geometry-based matching can be further improved by taking into account feature-based refinement. To achieve this, the descriptor maps are used, following:
@@
where $d(\cdot,\cdot)$ denotes the dot-product. As the channel number of $\mathbf{D}$ is large and the descriptor map lacks smoothness, a neighborhood searching (taking $\hat{\mathbf{u}}^{i}_{j}$ as initials), rather than optimization, is utilized to obtain $(\hat{\mathbf{u}}^{i}_{j})^{\prime}$.
To achieve sub-pixel data association, we asymmetrically upsample $\mathbf{D}^{ij}_{j}$ through bilinear interpolation:
@@
then perform the refinement around $(\hat{\mathbf{u}}^{i}_{j})^{\prime}$, following:
@@
III-B Visual Constraint based on Pointmap Alignment
Considering multiple image frames, we maintain both the pointmaps and their associated poses, represented as:
@@ III-B Visual Constraint based on Pointmap Alignment
is the camera-to-world Sim(3) transformation, $s\in\mathbb{R}^{+}$ denotes the scale, $\mathbf{R}\in\mathrm{SO}(3)$ denotes the rotation, and $\mathbf{t}\in\mathbb{R}^{3}$ denotes the translation. Essentially, $\mathbf{S}_{i}$ first scales the points in the camera frame to a unified scale and then transforms them into the world frame.
The self-referenced pointmap $\mathbf{X}_{i}$ is initialized using the two-view feed-forward results from Eq. (2), following:
@@
Specifically, assuming known camera intrinsics, the 3D points can be constrained to lie along their corresponding rays:
$\mathbf{X}_{i}\leftarrow\pi^{-1}\left(\mathbf{u}_{i},\left(\mathbf{X}_{i}\right)_{z}\right)$ (10)
where $\mathbf{u}_{i}$ are the grid-like pixel coordinates, $\left(\mathbf{X}_{i}\right)_{z}$ is the depth vector, $\pi(\cdot)$ is the camera projection model, whose inverse operation backprojects 2D pixels to 3D points. It is noted that, although feed-forward reconstruction is considered advantageous in that it obviates the requirement for camera parameters, in practice we found that camera distortions can introduce considerable instability. Therefore, we mainly consider the case where the camera parameters are known.
Figure 5: Illustration of different forms of visual constraints. Left: bundle adjustment, in which landmarks and camera poses are jointly optimized with little prior knowledge. Right: pointmap alignment, in which camera poses and pointmap scales are optimized, built upon the knowledge of the scaleless 3D structure.
Based on the dense association $\mathbf{u}^{i}_{j}$ produced by two-view matching, we can align two pointmaps and construct the following residuals (as shown in Fig. 4):
@@
where $\mathbf{S}^{i}_{j}=\mathbf{S}_{i}^{-1}\circ\mathbf{S}_{j}$ is the inter-frame relative Sim(3) transformation, $\circ$” denotes the transformation operation. The residuals in the upper part resemble the reprojection error with known depth and play a dominant role, whereas the residuals in the lower part are primarily used to handle pure rotation cases.
The inter-frame transformation can be directly estimated based on the above pointmap alignment residuals:
@@
where bi-directional projections are considered, $\mathbf{Q}_{ij}$, $\mathbf{Q}_{ji}$ are the weighting factors which consider network-outputted confidence and Huber robust kernel function. This optimization is used for pose tracking of a new frame.
It is noted that, the residual computation and the optimization process are dense due to the large number of pixels. Therefore, both the residuals and the Hessian information for each projection are computed on GPU. To be specific, the Hessian inforamtion is computed following:
@@
where $\bm{\eta}^{i}_{j}\in\mathbb{R}^{7}$ is the Lie algebra of $\mathbf{S}^{i}_{j}$, $\mathbf{J}^{r}_{ij}$ is the $\mathbf{r}_{ij}$-to-$\mathbf{S}^{i}_{j}$ Jacobian. After GPU processing, the Hessian-form inforamtion $(\mathbf{H}_{ij},\mathbf{v}_{ij})$ are in compact $(7,7)$ and $(7,1)$ shapes and can be processed efficiently on CPU.
When multiple frames within a neighborhood are considered, the pointmap alignment residuals can be constructed for all co-visible image pairs, thereby enabling joint estimation of multi-frame poses and scales:
@@
where $\mathcal{E}$ is the set of valid two-view projections, corresponding to a directed graph. During the optimization, relative-to-absolute Jacobians $\mathbf{J}^{ij}_{(i,j)}\in\mathbb{R}^{7\times 14}$ are utilized.
In contrast to BA commonly used in vision-based methods that jointly estimate point depths and camera poses, the above-described visual alignment model does not include point depths, and instead builds relatively independent constraints between pairwise images. This greatly simplifies the formulation of the problem, which is made possible by the strong capability of the visual foundation model to provide accurate estimates of 3D scene structure (up to a certain scale).
For comparison, we revisit the widely used BA technique and the re-prejection error formula, as shown below:
$\mathbf{r}_{ij}\left(\mathbf{T}^{i}_{j},\mathbf{z}_{j}\right)=\left[\begin{matrix}\mathbf{u}^{i}_{j}-\pi\left(\mathbf{T}^{i}_{j}\circ\pi^{-1}\left(\mathbf{u}_{j},\mathbf{z}_{j}\right)\right)\end{matrix}\right]$ (16)
$\left\{\hat{\mathbf{T}},\hat{\mathbf{z}}\right\}=\arg\min_{\left\{\mathbf{T},\mathbf{z}\right\}}\sum_{i,j\in\mathcal{K}}\left\|\mathbf{r}_{ij}\left(\mathbf{T}_{i},\mathbf{T}_{j},\mathbf{z}_{j}\right)\right\|^{2}_{\mathbf{Q}_{ij}}$ (17)
where $\mathbf{z}_{j}$ are the depths of landmarks in the source frame $j$, $\mathbf{u}_{j}$ are the pixel coordinates.
It can be seen that when BA constructs visual constraints, it relies on the parameterization of a large number of point depths $\mathbf{z}_{j}$; this greatly increases the scale of the problem and needs careful initialization for good convergence. For better efficiency, a two-step optimization can be adopted by eliminating landmarks first through Schur complement and then constructing pose constraints[38]. However, this process involves all frames related to $\mathbf{z}_{j}$, making the construction of visual constraints less flexible. These issues no longer exist in pointmap-alignment-based visual constraints, which can be easily modeled as pairwise residual terms.
It is noted that, this advantage relies on the assumption that the pointmap structure is accurate and scale is the only uncertain variable. Thanks to the spatial awareness of the feed-forward model, this assumption generally holds. Yet, under specific conditions, significant errors may be introduced, particularly when the points appear very far in image $j$ but very close in image $i$. In such cases, the depth uncertainty of the points in image $j$ will have a pronounced impact (corresponding to large disparities) in image $i$. This issue is not obvious in small-scale scenes, but becomes significant in large-scale outdoor environments, typically for the long-time forward-moving scenario.
To mitigate the influence of this uncertainty, we simply apply a mask to downweight the residuals in the above mentioned projection process, as shown below:
@@
IV Real-Time SLAM
Based on the above paradigm, visual alignment constraints operate on a frame-to-frame basis, involving only a very compact set of associated parameters. Beyond the shared common parameters, there exist no functional dependencies among distinct visual constraints. This characteristic enables highly flexible and modular construction of factor graphs. Nevertheless, several issues need to be addressed: